A rectangle has a perimeter of 34ft. and an area of 60 sq. ft. Find its dimensions.
Answer by umnikmo(75)

2(x+y)=34
x*y=60
x=60/y
60/y+y=17
y^2-17y+60=0
y=5
x=60/5=12